OpenStudy (anonymous):

Write the equation of a line in slope intercept form that is parallel to the line 2x-3y=12, and passes thought the point (2,6)? please help for fan!!

OpenStudy (triciaal):

Slope-intercept form for the equation of the line is y = mx + b parallel lines have the same slope 2x -3y =12 rearrange the equation to make y the subject y = 2/3(x) - 4 slope ,m = 2/3 parallel line passes through (2,6) x = 2, y = 6 and m = 2/3 plug the values and find b 6 = 2/3(2) + b ; b = 14/3 y = 2/3(x) + 14/3 this is the eqn of the line in slope-intercept form, parallel to the one given and passes through the point (2,6)

OpenStudy (anonymous):

thank u!!

OpenStudy (anonymous):

can u help with another one?

OpenStudy (triciaal):

you are welcome sure, go ahead and post it

OpenStudy (anonymous):

I meant to type perpendicular on my q, not parallel omg :(((((

OpenStudy (anonymous):

Write an equation of a line in slope-intercept form that is parallel to the line x+4y=6, and passes through the point (-8,5)???? is the other one

OpenStudy (triciaal):

it's ok perpendicular lines the product of the slope is -1 line 1 the slope is m then the slope of the perpendicular line line 2 is -1/m m*(-1/m) = -1

OpenStudy (triciaal):

follow the format and let me see what you have

OpenStudy (anonymous):

I thought you first had to put 2x-3y=12 into slope-intercept form, bit then you get a 'b' and the next step is to solve for be, that's where I was confused. I didn't know if there was a different rule for the problems like this given this way or not?

OpenStudy (triciaal):

y=-1/4(x) + 3/2 parallel or perpendicular will determine the slope of the 2nd line, m point given (-8, 5) so x = -8, y = 5 y = mx + b , find b write the new equation let me know if you need more help

OpenStudy (anonymous):

y=-1/4(x)+3/2 is what I got. but there is already a 'b' in that equation so I don't know how to solve for the next b, and what to do when i have 2 b's?

OpenStudy (triciaal):

The purpose of the first equation is to get m basically ignore everything else

OpenStudy (anonymous):

oh okay thank you so muc